# High-dimensional components of ${\rm SL}(2,\mathbb{C})$-character varieties of prime knots

### v1

@@ 3.1 A general structure
Note that $N(T_{0})=B(-(4n+1)/n)$, as illustrated in Figure 2, right. Given each generic $t\neq 0$, there exists $\rho_{0}\in\mathcal{R}_{t}^{\rm irr}(B(-\frac{4n+1}{n}))$, which is the same as some $\varrho_{0}\in\mathcal{R}_{t}(T_{0})$ with $\mathbf{g}_{\varrho_{0}}=\mathbf{e}$. Then $\mathbf{y}:=\varrho_{0}(y)=\varrho_{0}(T_{0}^{\rm ne})$ does not commute with $\mathbf{x}:=\varrho_{0}(x)=\varrho_{0}(T_{0}^{\rm se})$. This is because $\mathbf{y}=\mathbf{z}\mathbf{x}\mathbf{z}^{-1}$, where $\mathbf{z}=\varrho_{0}(z)$; if $\mathbf{x}\mathbf{y}=\mathbf{y}\mathbf{x}$, then $\mathbf{z}\mathbf{x}\mathbf{z}^{-1}\in\{\mathbf{x}^{\pm 1}\}$, so that $\mathbf{z}^{2}\mathbf{x}\mathbf{z}^{-2}=\mathbf{z}\mathbf{y}\mathbf{z}^{-1}=\mathbf{x}$, implying $\mathbf{y}\mathbf{z}=\mathbf{z}\mathbf{y}$, which contradicts the irreducibility of $\rho_{0}$.
@@ 3.2 Conditional structures
We partially answer Question 3.7, by showing that when
@@
for some $T_{i}\in\mathcal{T}_{2}^{2}$, there are two possibilities for $K$ to admit a HDC. In each case, the $T_{i}$’s are required to satisfy some conditions.
@@
  1. Case 1
    Suppose that for each generic $t$, there exist $\varrho_{i}\in\mathcal{R}_{t}(N(T_{i}))$, $i=1,2$ and $\varrho_{j}\in\mathcal{R}_{t}(D(T_{j}))$, $j=3,4$ such that ${\rm tr}_{v}(\varrho_{1})={\rm tr}_{v}(\varrho_{2})=a\in\mathbb{B}_{t}$, and ${\rm tr}_{h}(\varrho_{3})={\rm tr}_{h}(\varrho_{4})=b\in\mathbb{B}_{t}$. This requires $T_{1},T_{2}$ to be related in a certain way, and similarly for $T_{3},T_{4}$.
    Given $(\mathbf{x},\mathbf{y},\mathbf{z})\in G(t)^{3}$ with ${\rm tr}(\mathbf{x}\mathbf{y})=a$, ${\rm tr}(\mathbf{x}\mathbf{z})=b$, there exists $\rho\in\mathcal{R}_{t}(K)$ with $\rho(x_{1})=\rho(x_{2})=\rho(x_{3})=\rho(x_{4})=\mathbf{x}$, $\rho(y_{1})=\rho(y_{2})=\mathbf{y}$, $\rho(z_{1})=\rho(z_{2})=\mathbf{z}$. See the left part of Figure 3 for demonstration.
    The conjugacy classes of $(\mathbf{x},\mathbf{y},\mathbf{z})$ are parameterized by $c:={\rm tr}(\mathbf{y}\mathbf{z})$ and $r:={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(a,b,c;r)=0$. Therefore, $K$ admits a HDC.
  2. Case 2
    Suppose $T_{4}=\sigma(T_{3})$, where $\sigma$ is the reflection along the dotted line (see the right part of Figure 3), so $\mathsf{a}\mapsto\mathsf{a}^{\prime}:=\sigma(\mathsf{a})^{-1}$ establishes a bijection between directed arcs of $T_{3}$ and those of $T_{4}$. Furthermore, suppose that for each generic $t$, there exists $\varrho_{i}\in\mathcal{R}_{t}(N(T_{i}))$ with ${\rm tr}_{v}(\varrho_{i})=a_{i}\in\mathbb{B}_{t}$, $i=1,2$, and for generic $b$, there exists $\tau\in\mathcal{R}_{t}(T_{3})$ with ${\rm tr}_{v}(\varrho_{i})=b$.
    Given $(\mathbf{x},\mathbf{y},\mathbf{z})\in G(t)^{3}$ with ${\rm tr}(\mathbf{x}\mathbf{y})=a_{1}$, ${\rm tr}(\mathbf{y}\mathbf{z})=a_{2}$, and ${\rm tr}(\mathbf{x}\mathbf{z})$ being generic, by the assumption, there exists $\varrho\in\mathcal{R}(T_{3})$ with $\varrho(x_{1})=\mathbf{x}$, $\varrho(z_{1})=\mathbf{z}$. Put $\varrho^{\prime}(\mathsf{a}^{\prime})=\varrho(\mathsf{a})$ for each $\mathsf{a}^{\prime}$. Then $\varrho^{\prime}(\mathsf{c}^{\prime})=\varrho^{\prime}(\mathsf{a}^{\prime})\lrcorner\varrho^{\prime}(\mathsf{b}^{\prime})$ for any directed arcs $\mathsf{a}^{\prime},\mathsf{b}^{\prime},\mathsf{c}^{\prime}$ making a crossing of $T_{4}$ (see Figure 4), as $\varrho(\mathsf{c})=\varrho(\mathsf{a})\lrcorner\varrho(\mathsf{b})$. So $\varrho^{\prime}\in\mathcal{R}(T_{4})$, which satisfies $\varrho^{\prime}(x_{2})=\mathbf{x}$, $\varrho^{\prime}(z_{2})=\mathbf{z}$. Gluing $\varrho$ and $\varrho^{\prime}$ yields $\rho_{\mathbf{x}}^{\mathbf{z}}\in\mathcal{R}(T_{3}+T_{4})$ with $\rho_{\mathbf{x}}^{\mathbf{z}}(x_{i})=\mathbf{x}$ and $\rho_{\mathbf{x}}^{\mathbf{z}}(z_{i})=\mathbf{z}$ for $i=1,2$. By the assumption again, it can be extended to $\rho\in\mathcal{R}_{t}(K)$ with $\rho(x_{i})=\mathbf{x}$, $\rho(y_{i})=\mathbf{y}$, $\rho(z_{i})=\mathbf{z}$, $i=1,2$.
    The conjugacy classes of $(\mathbf{x},\mathbf{y},\mathbf{z})$ are parameterized by $b:={\rm tr}(\mathbf{x}\mathbf{z})$ and $r:={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(a_{1},a_{2},b;r)=0$. Thus, there is a surjection $\mathcal{X}_{t}^{\rm irr}(K)\to\{(b,r)\colon f_{t}(a_{1},a_{2},b;r)=0\}$. Therefore, $K$ admits a HDC.
@@ 4 Examples
In this section, we study three hyperbolic knots, explicitly describing the HDCs they admit. As a byproduct, we positively answer Question 1.1.
@@
$\gamma_{k}(r)=\begin{cases}(\mu^{k}-\mu^{-k})/(\mu-\mu^{-1}),&\mu\neq\pm 1\\ k\mu^{k-1},&\mu=\pm 1\end{cases}.$
@@
$\mathbf{x}^{k}=\gamma_{k}(r)\mathbf{x}-\gamma_{k-1}(r)\mathbf{e}.$
@@
If ${\rm tr}(\mathbf{a})={\rm tr}(\mathbf{b})$ and ${\rm tr}(\mathbf{a}\mathbf{b})=1$, then $\mathbf{a}\mathbf{b}\mathbf{a}^{k}\mathbf{b}\mathbf{a}=-\mathbf{b}^{k-2}$ for each $k\in\mathbb{Z}$; in particular, $\mathbf{a}\mathbf{b}^{2}\mathbf{a}=-\mathbf{b}^{-2}$.
@@
$\mathbf{a}\mathbf{b}\mathbf{a}^{k}\mathbf{b}\mathbf{a}\mathbf{b}^{2}$ $=\mathbf{a}\mathbf{b}(\gamma_{k}(t)\mathbf{a}-\gamma_{k-1}(t)\mathbf{e})\mathbf{b}\mathbf{a}\mathbf{b}^{2}$
$=\gamma_{k}(t)(\mathbf{a}\mathbf{b})^{3}\cdot\mathbf{b}-\gamma_{k-1}(t)\mathbf{a}\mathbf{b}^{2}\mathbf{a}\cdot\mathbf{b}^{2}$
$=-\gamma_{k}(t)\mathbf{b}+\gamma_{k-1}(t)\mathbf{e}=-\mathbf{b}^{k}.$
@@ 4.1 The $(3,3,3,4)$-pretzel knot
Let $P$ denote the $(3,3,3,4)$-pretzel knot, as shown in Figure 5, left.
@@
determines a representation $\rho_{\underline{\mathbf{x}}}\in\mathcal{R}_{t}^{\rm irr}(K)$ such that $\mathbf{x}_{i}=\rho_{\underline{\mathbf{x}}}(x_{i})$, $1\leq i\leq 4$ and $\mathbf{g}_{\rho_{\underline{\mathbf{x}}}}=\mathbf{e}$.
@@
is injective, and its image consists of $(t_{13},t_{123},t_{134})$ with $f_{t}(1,1,t_{13};t_{123})=f_{t}(1,0,t_{13};t_{134})=0$, $t_{13}\neq t^{2}-2$, and $t_{13}\neq 2$ if $t^{2}\in\{2,3\}$.
@@
Given any $t_{13},t_{123}$ with $f_{t}(1,1,t_{13};t_{123})=0$ and $t_{13}\in\mathbb{B}_{t}$, up to conjugacy there uniquely exists $(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})\in G(t)^{3}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3})=t_{13}$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3})=t_{123}$. Since $t_{13}\neq 2,t^{2}-2$, by Lemma 2.1 (ii-2), any $t_{134}$ with $f_{t}(1,0,t_{13};t_{134})=0$ further determines a unique $\mathbf{x}_{4}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})=0$, ${\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3}\mathbf{x}_{4})=t_{134}$.
Now assume $t_{13}\in\{2,t^{2}-2\}$. Given any $t_{134}$ with $f_{t}(1,0,t_{13};t_{134})=0$, up to conjugacy there is a unique $(\mathbf{x}_{1},\mathbf{x}_{3},\mathbf{x}_{4})\in G(t)^{3}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})=0$, ${\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3}\mathbf{x}_{4})=t_{134}$. Since ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})\neq{\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})$, we have $\mathbf{x}_{1}\neq\mathbf{x}_{3}$. Note that $\mathbf{x}_{1}=\mathbf{x}_{3}^{-1}$ if and only if $t^{2}-2={\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})+{\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1$, i.e. $t^{2}=3$; when this holds, there does not exist $\mathbf{x}_{2}$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$, as ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})+{\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})\neq t^{2}-2$. Conversely, suppose $(t^{2},t_{13})\neq(3,2)$, then $(\mathbf{x}_{1},\mathbf{x}_{3})$ is reducible and $\mathbf{x}_{1}\mathbf{x}_{3}\neq\mathbf{x}_{3}\mathbf{x}_{1}$. Following the proof of Lemma 3.2, in the context of Point 2 there, we can deduce $\varepsilon=-1$, so that $t_{13}=2$, and $t^{2}\neq 2$. When these hold, there exists $\mathbf{x}_{2}$ satisfying ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$; it is fixed by $t_{123}={\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3})$ which satisfies $f_{t}(1,1,2;t_{123})=0$. ∎
Consequently, $P$ admits a HDC isomorphic to
$\big\{(t,t_{13},t_{123},t_{134})\colon$ $t_{13}\neq t^{2}-2\ \text{and\ }(t^{2},t_{13})\neq(2,2),(3,2),$

### v2

@@ 3.1 A general structure
Note that $N(T_{0})=B(-(4n+1)/n)$, as illustrated in Figure 2, right. For each generic $t$, there exists $\tau\in\mathcal{R}_{t}^{\rm irr}(N(T_{0}))$, which is the same as some $\varrho_{0}\in\mathcal{R}_{t}(T_{0})$ with $\mathbf{g}_{\varrho_{0}}=\mathbf{e}$. Then $\mathbf{y}:=\tau(y)=\varrho_{0}(T_{0}^{\rm ne})$ does not commute with $\mathbf{x}:=\tau(x)=\varrho_{0}(T_{0}^{\rm se})$. This is because $\mathbf{y}=\mathbf{z}\lrcorner\mathbf{x}^{-1}$, where $\mathbf{z}=\tau(z)$; if $\mathbf{x}\mathbf{y}=\mathbf{y}\mathbf{x}$, then $\mathbf{z}\lrcorner\mathbf{x}=\mathbf{x}^{\varepsilon}$ for some $\varepsilon\in\{\pm 1\}$, so that
$\mathbf{z}^{2}\lrcorner\mathbf{x}=\mathbf{z}\lrcorner(\mathbf{z}\lrcorner\mathbf{x})=\mathbf{z}\lrcorner\mathbf{x}^{\varepsilon}=(\mathbf{z}\lrcorner\mathbf{x})^{\varepsilon}=(\mathbf{x}^{\varepsilon})^{\varepsilon}=\mathbf{x},$
implying $\mathbf{x}\in\{\mathbf{z}^{\pm 2}\}$, so $\mathbf{y}\mathbf{z}=\mathbf{z}\mathbf{y}$, which contradicts the irreducibility of $\tau$.
@@ 3.2 Conditional structures
We partially answer Question 3.6, by showing that when
@@
for some $T_{i}\in\mathcal{T}_{2}^{2}$, there are two possibilities for $K$ to admit a HDC. In either case, the $T_{i}$’s are required to satisfy some conditions.
@@
Case 1. Suppose that for each generic $t$, there exist $\varrho_{i}\in\mathcal{R}_{t}(N(T_{i}))$, $i=1,2$ and $\varrho_{j}\in\mathcal{R}_{t}(D(T_{j}))$, $j=3,4$ such that ${\rm tr}_{v}(\varrho_{1})={\rm tr}_{v}(\varrho_{2})=:a\in\mathbb{B}_{t}$, and ${\rm tr}_{h}(\varrho_{3})={\rm tr}_{h}(\varrho_{4})=:b\in\mathbb{B}_{t}$. This requires $T_{1},T_{2}$ to be related in a certain way, and similarly for $T_{3},T_{4}$.
Given $(\mathbf{x},\mathbf{y},\mathbf{z})\in G(t)^{3}$ with ${\rm tr}(\mathbf{x}\mathbf{y})=a$, ${\rm tr}(\mathbf{x}\mathbf{z})=b$, similarly as in the proof of Theorem 3.2, there exist $\rho_{i}\in\mathcal{R}_{t}(T_{i})$, $1\leq i\leq 4$ such that
$\rho_{1}(T_{1}^{\rm nw})$ $=\mathbf{x},\qquad$ $\rho_{1}(T_{1}^{\rm ne})$ $=\mathbf{x}^{-1},\qquad$ $\rho_{1}(T_{1}^{\rm sw})$ $=\mathbf{y},\qquad$ $\rho_{1}(T_{1}^{\rm ne})$ $=\mathbf{y}^{-1},$
$\rho_{2}(T_{2}^{\rm se})$ $=\mathbf{x},\qquad$ $\rho_{2}(T_{2}^{\rm sw})$ $=\mathbf{x}^{-1},\qquad$ $\rho_{2}(T_{2}^{\rm ne})$ $=\mathbf{y},\qquad$ $\rho_{2}(T_{2}^{\rm nw})$ $=\mathbf{y}^{-1},$
$\rho_{3}(T_{3}^{\rm nw})$ $=\mathbf{x},\qquad$ $\rho_{3}(T_{3}^{\rm sw})$ $=\mathbf{x}^{-1},\qquad$ $\rho_{3}(T_{3}^{\rm ne})$ $=\mathbf{z},\qquad$ $\rho_{3}(T_{3}^{\rm se})$ $=\mathbf{z}^{-1},$
$\rho_{4}(T_{4}^{\rm se})$ $=\mathbf{x},\qquad$ $\rho_{4}(T_{4}^{\rm ne})$ $=\mathbf{x}^{-1},\qquad$ $\rho_{4}(T_{4}^{\rm sw})$ $=\mathbf{z},\qquad$ $\rho_{4}(T_{4}^{\rm nw})$ $=\mathbf{z}^{-1}.$
We can glue the $\rho_{i}$’s into some $\rho\in\mathcal{R}_{t}^{\rm irr}(K)$ with $\rho(x_{i})=\mathbf{x}$ for $1\leq i\leq 4$ and $\rho(y_{i})=\mathbf{y}$, $\rho(z_{i})=\mathbf{z}$ for $i=1,2$. See the left part of Figure 3.
By Lemma 2.3, the conjugacy classes of $(\mathbf{x},\mathbf{y},\mathbf{z})$ are parameterized by $c:={\rm tr}(\mathbf{y}\mathbf{z})$ and $r:={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(a,b,c;r)=0$. Therefore, $K$ admits a HDC.
@@
Case 2. Suppose $T_{4}=\sigma(T_{3})$, where $\sigma$ is the reflection along the dotted line (see the right part of Figure 3), so $\mathsf{a}\mapsto\mathsf{a}^{\prime}:=\sigma(\mathsf{a})^{-1}$ establishes a bijection between directed arcs of $T_{3}$ and those of $T_{4}$. Furthermore, suppose that for each generic $t$, there exists $\varrho_{i}\in\mathcal{R}_{t}(N(T_{i}))$ with ${\rm tr}_{v}(\varrho_{i})=a_{i}\in\mathbb{B}_{t}$, $i=1,2$, and for each generic $b$, there exists $\tau\in\mathcal{R}_{t}(T_{3})$ with ${\rm tr}_{v}(\tau)=b$.
Given $(\mathbf{x},\mathbf{y},\mathbf{z})\in G(t)^{3}$ with ${\rm tr}(\mathbf{x}\mathbf{y})=a_{1}$, ${\rm tr}(\mathbf{y}\mathbf{z})=a_{2}$ and ${\rm tr}(\mathbf{x}\mathbf{z})$ being generic, there exist $\rho_{i}\in\mathcal{R}_{t}(T_{i})$, $i=1,2,3$ such that
$\rho_{1}(T_{1}^{\rm nw})$ $=\mathbf{x},\qquad$ $\rho_{1}(T_{1}^{\rm ne})$ $=\mathbf{x}^{-1},\qquad$ $\rho_{1}(T_{1}^{\rm sw})$ $=\mathbf{y},\qquad$ $\rho_{1}(T_{1}^{\rm ne})$ $=\mathbf{y}^{-1},$
$\rho_{2}(T_{2}^{\rm ne})$ $=\mathbf{y},\qquad$ $\rho_{2}(T_{2}^{\rm nw})$ $=\mathbf{y}^{-1},\qquad$ $\rho_{2}(T_{2}^{\rm se})$ $=\mathbf{z},\qquad$ $\rho_{2}(T_{2}^{\rm sw})$ $=\mathbf{z}^{-1},$
$\rho_{3}(T_{3}^{\rm nw})$ $=\mathbf{z},\qquad$ $\rho_{3}(T_{3}^{\rm sw})$ $=\mathbf{x}^{-1}.$
Define $\rho_{4}\in\mathcal{R}(T_{4})$ by setting $\rho_{4}(\mathsf{a}^{\prime})=\rho_{3}(\mathsf{a})$ for each directed arc $\mathsf{a}^{\prime}$ of $T_{4}$. Then $\rho_{4}(\mathsf{c}^{\prime})=\rho_{4}(\mathsf{a}^{\prime})\lrcorner\rho_{4}(\mathsf{b}^{\prime})$ for any directed arcs $\mathsf{a}^{\prime},\mathsf{b}^{\prime},\mathsf{c}^{\prime}$ making a crossing in $T_{4}$ (see Figure 4), as $\rho_{3}(\mathsf{c})=\rho_{3}(\mathsf{a})\lrcorner\rho_{3}(\mathsf{b})$. So indeed $\rho_{4}\in\mathcal{R}(T_{4})$. It satisfies $\rho_{4}(T_{4}^{\rm ne})=\mathbf{z}^{-1}$, $\rho_{4}(T_{4}^{\rm se})=\mathbf{x}$. We can glue the $\rho_{i}$’s into some $\rho\in\mathcal{R}_{t}^{\rm irr}(K)$ with $\rho(x_{i})=\mathbf{x}$, $\rho(y_{i})=\mathbf{y}$, $\rho(z_{i})=\mathbf{z}$ for $i=1,2$.
The conjugacy classes of $(\mathbf{x},\mathbf{y},\mathbf{z})$ are parameterized by $b:={\rm tr}(\mathbf{x}\mathbf{z})$ and $r:={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(a_{1},a_{2},b;r)=0$. Therefore, $K$ admits a HDC.
@@ 4 Examples
In this section, we study three hyperbolic knots, explicitly describing the HDCs they admit. Then we positively answer Question 1.1.
We clarify some basic facts first.
@@
$\omega_{k}(r)=\begin{cases}(\mu^{k}-\mu^{-k})/(\mu-\mu^{-1}),&\mu\neq\pm 1\\ k\mu^{k-1},&\mu=\pm 1\end{cases}.$
@@
$\mathbf{x}^{k}=\omega_{k}(r)\mathbf{x}-\omega_{k-1}(r)\mathbf{e}.$
@@
If $\mathbf{a},\mathbf{b}\in G(t)$ and ${\rm tr}(\mathbf{a}\mathbf{b})=1$, then $\mathbf{a}\mathbf{b}\mathbf{a}^{k}\mathbf{b}\mathbf{a}=-\mathbf{b}^{k-2}$ for each $k\in\mathbb{Z}$; in particular, $\mathbf{a}\mathbf{b}^{2}\mathbf{a}=-\mathbf{b}^{-2}$.
@@
$\mathbf{a}\mathbf{b}\mathbf{a}^{k}\mathbf{b}\mathbf{a}\mathbf{b}^{2}$ $=\mathbf{a}\mathbf{b}(\omega_{k}(t)\mathbf{a}-\omega_{k-1}(t)\mathbf{e})\mathbf{b}\mathbf{a}\mathbf{b}^{2}$
$=\omega_{k}(t)(\mathbf{a}\mathbf{b})^{3}\cdot\mathbf{b}-\omega_{k-1}(t)\mathbf{a}\mathbf{b}^{2}\mathbf{a}\cdot\mathbf{b}^{2}$
$=-\omega_{k}(t)\mathbf{b}+\omega_{k-1}(t)\mathbf{e}=-\mathbf{b}^{k},$
@@
Figure 5:
Consider the tangle $[n]$ with $n>0$, as shown in Figure 5.
Suppose $\tau\in\mathcal{R}_{t}([n])$, with $\tau(a_{i})=\mathbf{a}_{i}$, and ${\rm tr}(\mathbf{a}_{0}\mathbf{a}_{1})=s$. Then $\mathbf{a}_{k+1}\mathbf{a}_{k}=\mathbf{a}_{1}\mathbf{a}_{0}$ for $1\leq k\leq n$, hence $\mathbf{a}_{k+2}=(\mathbf{a}_{k+1}\mathbf{a}_{k})\lrcorner\mathbf{a}_{k}=(\mathbf{a}_{1}\mathbf{a}_{0})\lrcorner\mathbf{a}_{k}$. Consequently,
$\mathbf{a}_{3}=\mathbf{a}_{1}\mathbf{a}_{0}\mathbf{a}_{1}(\mathbf{a}_{1}\mathbf{a}_{0})^{-1},\qquad\mathbf{a}_{4}=(\mathbf{a}_{1}\mathbf{a}_{0})^{2}\lrcorner\mathbf{a}_{0}.$
As special cases, if $s=1$, then $(\mathbf{a}_{1}\mathbf{a}_{0})^{3}=-\mathbf{e}$, so
$\mathbf{a}_{3}=-\mathbf{a}_{1}\mathbf{a}_{0}\mathbf{a}_{1}(\mathbf{a}_{1}\mathbf{a}_{0})^{2}=-\mathbf{a}_{1}\mathbf{a}_{0}\mathbf{a}_{1}^{2}\mathbf{a}_{0}\mathbf{a}_{1}\cdot\mathbf{a}_{0}=\mathbf{a}_{0};$
if $s=0$, then $(\mathbf{a}_{1}\mathbf{a}_{0})^{2}=-\mathbf{e}$, hence $\mathbf{a}_{4}=\mathbf{a}_{0}$.
By the formulas in [3, Example 3.4], with $h=1$, $r=t^{2}-s$,
${\rm tr}(\mathbf{a}_{1}^{-1}\mathbf{a}_{3})$ $=2+\grave{z}=2+(s+2-t^{2})(s-2),$ (1)
${\rm tr}(\mathbf{a}_{0}^{-1}\mathbf{a}_{3})$ $=2+z=2+(t^{2}-s-2)(s-1)^{2}.$ (2)
@@ 4.1 The $(3,3,3,4)$-pretzel knot
Let $P$ denote the $(3,3,3,4)$-pretzel knot, as shown in Figure 6.
@@
gives rise to a unique $\rho_{\underline{\mathbf{x}}}\in\mathcal{R}_{t}^{\rm irr}(K)$ with $\mathbf{x}_{i}=\rho_{\underline{\mathbf{x}}}(x_{i})$, $1\leq i\leq 4$ and $\mathbf{g}_{\rho_{\underline{\mathbf{x}}}}=\mathbf{e}$.
@@
is injective, and its image consists of $(t_{13},t_{123},t_{134})$ with $f_{t}(1,1,t_{13};t_{123})=f_{t}(1,0,t_{13};t_{134})=0$, $t_{13}\neq t^{2}-2$, and $t_{13}\neq 2$ if $t^{2}\in\{1,2\}$.
@@
Given any $t_{13}\in\mathbb{B}_{t}$ and $t_{123}$ with $f_{t}(1,1,t_{13};t_{123})=0$, up to conjugacy there exists a unique $(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})\in G(t)^{3}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3})=t_{13}$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3})=t_{123}$. Since $t_{13}\in\mathbb{B}_{t}$, by Lemma 2.3 (ii-2), any $t_{134}$ with $f_{t}(1,0,t_{13};t_{134})=0$ further determines a unique $\mathbf{x}_{4}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})=0$, ${\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3}\mathbf{x}_{4})=t_{134}$.
From now on, let $t_{13}\in\{2,t^{2}-2\}$. Given any $t_{134}$ with $f_{t}(1,0,t_{13};t_{134})=0$, up to conjugacy there exits a unique $(\mathbf{x}_{1},\mathbf{x}_{3},\mathbf{x}_{4})\in G(t)^{3}$ such that
${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})=0,\qquad{\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1,\qquad{\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3}\mathbf{x}_{4})=t_{134}.$
Since ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})\neq{\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})$, we have $\mathbf{x}_{1}\neq\mathbf{x}_{3}$.
  • If $\mathbf{x}_{1}=\mathbf{x}_{3}^{-1}$, then $t_{13}=2$, and $\mathbf{x}_{1}+\mathbf{x}_{3}=t\mathbf{e}$, so
    $1={\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})+{\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})={\rm tr}((\mathbf{x}_{1}+\mathbf{x}_{3})\mathbf{x}_{4})=t\cdot{\rm tr}(\mathbf{x}_{4})=t^{2}.$
  • Conversely, if $(t^{2},t_{13})=(1,2)$, then $f_{t}(1,0,t_{13};t_{134})=0$ forces $t_{134}=t$; since $(\mathbf{x}_{3},\mathbf{x}_{4})$ is irreducible and now ${\rm tr}(\mathbf{x}_{1}\mathbf{c})={\rm tr}(\mathbf{x}_{3}^{-1}\mathbf{c})$ for $\mathbf{c}\in\{\mathbf{x}_{3},\mathbf{x}_{4},\mathbf{x}_{3}\mathbf{x}_{4}\}$, by the uniqueness in Lemma 2.3 (ii-2), $\mathbf{x}_{1}=\mathbf{x}_{3}^{-1}$.
Hence $(t^{2},t_{13})=(1,2)$ if and only if $\mathbf{x}_{1}=\mathbf{x}_{3}^{-1}$. When $(t^{2},t_{13})=(1,2)$, there does not exist $\mathbf{x}_{2}$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$: otherwise, we would deduce ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})+{\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=2\neq t^{2}={\rm tr}((\mathbf{x}_{1}+\mathbf{x}_{3})\mathbf{x}_{2})$.
Now suppose $(t^{2},t_{13})\neq(1,2)$. Then $(\mathbf{x}_{1},\mathbf{x}_{3})$ is reducible and $\mathbf{x}_{1}\mathbf{x}_{3}\neq\mathbf{x}_{3}\mathbf{x}_{1}$. Following the proof of Lemma 3.1, in the context of Point 2 there, we can deduce $\varepsilon=-1$, so that $t_{13}=2$, and $t^{2}\neq 2$. When these hold and given $t_{123}$ with $f_{t}(1,1,2;t_{123})=0$, there exists a unique $\mathbf{x}_{2}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$ and ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3})=t_{123}$. ∎
In conclusion, $P$ admits a HDC isomorphic to
$\big\{(t,t_{13},t_{123},t_{134})\colon$ $t_{13}\neq t^{2}-2\ \text{and\ }(t^{2},t_{13})\neq(2,2),(1,2),$